OpenStudy (anonymous):

Find the smallest positive value of the variable so that each of the following represents an integer. Show a full solution of how it is done please.

OpenStudy (anonymous):

a) \[\sqrt{3p-1}\] b) \[\sqrt{\frac{ y+5 }{ 2 }}\]

OpenStudy (mathstudent55):

a) \( \sqrt{3p - 1} = 0 \) \(3p - 1 = 0 \) \(3p = 1 \) \( p = \dfrac{1}{3} \)

OpenStudy (mathstudent55):

b) \( \sqrt{ \dfrac{y + 5}{2} } = 0 \) \( \dfrac{y + 5}{2} = 0 \) \( y + 5 = 0 \) \(y = -5\)

OpenStudy (anonymous):

Quick question: how did you get rid of the square root?

OpenStudy (mathstudent55):

Square both sides.